Stress at Work: in a poll conducted by the General Social Survey, 81% of respondents said that their jobs were sometimes or always stressful. 200 workers are chosen at random.
a. approximate the probability that 160 or fewer find their jobs stressful. please show work to explain answer
Answer #1

Answer: Given that h= 200 2=160 P 160 - 16.00 200 P=0.81 0.8-0.81 PCÊ 50.8)= e(P-P 200 (PCI-D 6.81(1-0-81) In 200 = P(25-0.36